### v1

Abstract.
Let $G$ be a simple graph on $n$ vertices with chromatic number $\chi$, and let $\lambda_{n}$ denote the least adjacency eigenvalue. Solving a conjecture of Fan, Yu and Wang [Electron. J. Combin., 2012], we prove that when $3\leq\chi\leq n-1$, the chromatic number satisfies the following upper bound:
$\chi\leq\left(\frac{n}{2}+1+\lambda_{n}\right)+\sqrt{\left(\frac{n}{2}+1+\lambda_{n}\right)^{2}-4(\lambda_{n}+1)\left(\lambda_{n}+\frac{n}{2}\right)},$
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\tfrac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even. This extends the validity of the Fan–Yu–Wang bound from the range $3\leq\chi\leq\frac{n}{2}$ to the full range $3\leq\chi\leq n-1$.
We also compare this bound with the well known bound due to Wilf that $\chi\leq 1+\lambda_{1}$, where $\lambda_{1}$ denotes the largest eigenvalue. In particular we show that while Wilf’s bound is an upper bound for some parameters larger than $\chi$, this bound using $\lambda_{n}$ is not an upper bound for these parameters.
Key words and phrases:
Chromatic number; Least adjacency eigenvalue; Spectral graph theory
2020 Mathematics Subject Classification:
Primary 05C50.
1. Introduction
Let $G=(V,E)$ be a simple graph of order $n$ with vertex set $V=V(G)$ and edge set $E=E(G)$. The adjacency matrix of $G$ is the $n\times n$ 011 matrix $A(G)=[a_{ij}]$, where $a_{ij}=1$ if $v_{i}$ is adjacent to $v_{j}$ and $a_{ij}=0$ otherwise. The eigenvalues of $A(G)$ are called the eigenvalues of the graph $G$. Since $A(G)$ is symmetric, all eigenvalues are real; we list them as
$\lambda_{1}(G)\geq\lambda_{2}(G)\geq\cdots\geq\lambda_{n}(G).$
The chromatic number of a graph $G$, denoted by $\chi(G)$, is the least number of colors needed to color $V(G)$ so that adjacent vertices receive distinct colors. There has been extensive research on the connection between eigenvalues and the chromatic number of a graph. Wilf [6] proved that $\chi(G)\leq 1+\lambda_{1}(G)$, with equality if and only if $G$ is a complete graph or an odd cycle. Hoffman [4] showed that $\chi(G)\geq 1-\frac{\lambda_{1}(G)}{\lambda_{n}(G)}$. Elphick, Tang and Zhang [2] proved lower bounds on chromatic numbers in terms of $p$-energies, which substantially generalize the Hoffman bound.
We use the following notation. Let $K_{n}$ (resp. $O_{n}$) denote the complete (resp. empty) graph on $n$ vertices; note that $O_{n}=nK_{1}$, the disjoint union of $n$ isolated vertices. For two vertex-disjoint graphs $H_{1}=(V_{1},E_{1})$ and $H_{2}=(V_{2},E_{2})$, their disjoint union is
$H_{1}\cup H_{2}\quad\text{with}\quad V(H_{1}\cup H_{2})=V_{1}\cup V_{2},\ \ E(H_{1}\cup H_{2})=E_{1}\cup E_{2},$
and their join is
$H_{1}\vee H_{2}\quad\text{with}\quad V(H_{1}\vee H_{2})=V_{1}\cup V_{2},\ \ E(H_{1}\vee H_{2})=E_{1}\cup E_{2}\cup\{uv:\ u\in V_{1},\ v\in V_{2}\}.$
(Equivalently, every vertex of $H_{1}$ is made adjacent to every vertex of $H_{2}$.) We also write $K_{a,b}$ for the complete bipartite graph with parts of sizes $a$ and $b$; note that $K_{a,b}=O_{a}\vee O_{b}$.
It is immediate that $\chi(G)=n$ if and only if $G=K_{n}$; $\chi(G)=1$ if and only if $G=O_{n}$. Moreover, $\chi(G)=2$ precisely when $G$ is bipartite. A result of Constantine [1] (see also [3]) shows that among all $n$-vertex graphs (or all $n$-vertex bipartite graphs), the unique graph minimizing the least eigenvalue is the complete bipartite graph $K_{\lceil n/2\rceil,\lfloor n/2\rfloor}=O_{\lceil n/2\rceil}\vee O_{\lfloor n/2\rfloor}$. Accordingly, the present work focuses on graphs with $\chi(G)\geq 3$.
In [3, Theorems 3.1 and 3.2], Fan, Yu and Wang established an upper bound for the chromatic number of a graph in terms of its order and least adjacency eigenvalue, which also appears in Stanić’s book [5, Theorem 3.11].
Theorem 1.1 ([3]).
Let $G$ be a graph of order $n$ with chromatic number $3\leq\chi\leq\frac{n}{2}$ and least adjacency eigenvalue $\lambda_{n}$. Then
$\chi\leq\left(\frac{n}{2}+1+\lambda_{n}\right)+\sqrt{\left(\frac{n}{2}+1+\lambda_{n}\right)^{2}-4(\lambda_{n}+1)\left(\lambda_{n}+\frac{n}{2}\right)},$
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\frac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\frac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
Fan, Yu and Wang [3] conjectured that the above bound continues to hold for $\frac{n}{2}<\chi\leq n-1$. In this paper we resolve this conjecture. This is achieved by replacing the calculus-based approach used in [3, Section 3] with a purely algebraic approach.
2. Preliminaries
We follow the notation in [3]. Given positive integers $n$ and $\chi$ with $3\leq\chi\leq n-1$, we introduce a family of graphs defined by
$G(a,a_{0}):=\left(K_{a}\cup O_{a_{0}}\right)\vee\left(K_{b}\cup O_{b_{0}}\right),$
where
$1\leq a\leq\chi-1,\qquad b=\chi-a,\qquad a_{0},b_{0}\geq 0,\qquad a_{0}+b_{0}=n-\chi.$ (2.1)
We call a pair $(a,a_{0})$ feasible if the integers $a$ and $a_{0}$ satisfy (2.1). It is immediate that $\chi(G(a,a_{0}))=\chi$, and since $G(a,a_{0})$ is connected but not complete, we have $\lambda_{n}(G(a,a_{0}))<-1$.
By [3, Eq. (2.2)], the least eigenvalue $\lambda_{n}(G(a,a_{0}))$ is the smallest real root of the quartic polynomial:
$f(a,a_{0},\lambda):=\lambda^{2}(\lambda-a+1)(\lambda-b+1)-\bigl[(b+b_{0})\lambda-b_{0}(b-1)\bigr]\bigl[(a+a_{0})\lambda-a_{0}(a-1)\bigr],$ (2.2)
where $b=\chi-a$ and $b_{0}=n-\chi-a_{0}$, and throughout the paper $n$ and $\chi$ are regarded as fixed numbers.
We recall the following structural lemma from [3, Lemma 2.1].
Lemma 2.1 ([3]).
Among all graphs of order $n$ and chromatic number $\chi$, where $3\leq\chi\leq n-1$, if a graph $G$ is one whose least eigenvalue attains the minimum, then $G$ must be of the form $G(a,a_{0})$ for some $a$ and $a_{0}$.
We also need the following lemma concerning the number of negative eigenvalues of the polynomial $f(a,a_{0},\lambda)$ defined in (2.2).
Lemma 2.2.
For every feasible pair $(a,a_{0})$, the polynomial $f(a,a_{0},\lambda)$ has exactly one negative real root (counted with multiplicity). Moreover, $f(a,a_{0},0)\leq 0$ and $\lim_{\lambda\to-\infty}f(a,a_{0},\lambda)=+\infty$.
Proof.
Write $g(x):=f(a,a_{0},-x)$. Let
$A:=a-1\geq 0,\quad B:=b-1\geq 0,\quad c_{1}:=b+b_{0},\quad d_{1}:=b_{0}(b-1)=b_{0}B,$
$c_{2}:=a+a_{0},\quad d_{2}:=a_{0}(a-1)=a_{0}A,$
where $b=\chi-a$ and $b_{0}=n-\chi-a_{0}$. A direct expansion gives
$g(x)$ $=x^{2}(x+A)(x+B)-(c_{1}x+d_{1})(c_{2}x+d_{2})$
$=x^{4}+(a+b-2)x^{3}+\left((a-1)(b-1)-(a+a_{0})(b+b_{0})\right)x^{2}$
$\qquad-\left((b+b_{0})a_{0}(a-1)+(a+a_{0})b_{0}(b-1)\right)x-a_{0}(a-1)b_{0}(b-1).$
Hence the coefficient signs of $g$ are
+,0,,0,0.+\ $\quad\geq$ 0\ $\quad-\ ,\quad\leq$ 0\ $\quad\leq$ 0.
Ignoring possible zeros, the sign sequence has exactly one sign change (from ++ to -). By Descartes’ rule of signs, $g$ has exactly one positive real root with multiplicity $1$. Therefore $f(a,a_{0},\lambda)$ has exactly one real root (counted with multiplicity) in $(-\infty,0)$.
Finally, $f(a,a_{0},0)=-d_{1}d_{2}=-b_{0}(b-1)\,a_{0}(a-1)\leq 0$, and since the leading coefficient of $f$ is +1+1, we have $\lim_{\lambda\to-\infty}f(a,a_{0},\lambda)=+\infty$. ∎
3. Main results
In this section we extend [3, Theorems 3.1 and 3.2] to the full range $3\leq\chi\leq n-1$.
Theorem 3.1.
Let $G$ be a simple graph of order $n$ with chromatic number $\chi$ satisfying $3\leq\chi\leq n-1$. Then the least adjacency eigenvalue $\lambda_{n}$ satisfies
$\lambda_{n}\geq-\frac{n-\chi+2+\sqrt{(n-\chi-2)^{2}+4\chi(n-\chi)}}{4},$
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\frac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\frac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
Proof.
By Lemma 2.1, to determine the extremal value of $\lambda_{n}(G)$ among all graphs of order $n$ and chromatic number $\chi$, it suffices to minimize the smallest real root of $f(a,a_{0},\lambda)$ over all feasible pairs $(a,a_{0})$. Our goal is to prove that, for every feasible pair $(a,a_{0})$,
[…]
Let $\xi$ denote the smallest real root of $f\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2},\lambda\right)$. In view of Lemma 2.2, each $f(a,a_{0},\lambda)$ has exactly one real root on $(-\infty,0)$; hence Eq. (3.1) is equivalent to showing that
$f(a,a_{0},\xi)\ \geq\ 0\qquad\text{for every feasible pair }(a,a_{0}).$ (3.2)
Indeed, if $f(a,a_{0},\xi)\geq 0$, then $\xi$ lies to the left of (or coincides with) the unique negative root of $f(a,a_{0},\lambda)$, so the smallest real root of $f(a,a_{0},\lambda)$ is at least $\xi$.
Fix $n,\chi$ and write
$p:=\frac{n-\chi}{2}>0,\quad q:=\frac{\chi}{2}-1>0,\quad M:=\frac{n}{2}\lambda,$ (3.3)
$a=\frac{\chi}{2}+t,\quad b=\frac{\chi}{2}-t,\quad a_{0}=p+s,\quad b_{0}=p-s,$ (3.4)
so that $|t|\leq q$ and $|s|\leq p$. Define
$\Delta(s,t;\lambda):=f(a,a_{0},\lambda)-f\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2},\lambda\right).$
The next claim is a purely algebraic identity.
Claim 1.
For all $\lambda\in\mathbb{R}$,
$\Delta(s,t;\lambda)=p(p-2\lambda)t^{2}+(\lambda-q)^{2}s^{2}-s^{2}t^{2}+2\lambda(\lambda+1)st.$
Proof of Claim 1.
Write
$L:=(b+b_{0})\lambda-b_{0}(b-1),\qquad R:=(a+a_{0})\lambda-a_{0}(a-1),$
a direct calculation with (3.3) and (3.4) gives
$L=(M-pq)-st-\left(s(\lambda-q)+t(\lambda-p)\right),\quad R=(M-pq)-st+\left(s(\lambda-q)+t(\lambda-p)\right),$
hence
$LR=(M-pq)^{2}-2(M-pq)st-\left[(\lambda-q)^{2}s^{2}+(\lambda-p)^{2}t^{2}+2(\lambda-q)(\lambda-p)st\right]+s^{2}t^{2}.$
Using
$(\lambda-a+1)(\lambda-b+1)=(\lambda-q-t)(\lambda-q+t)=(\lambda-q)^{2}-t^{2},$
we know that
$f(a,a_{0},\lambda)=\lambda^{2}\left[(\lambda-q)^{2}-t^{2}\right]-LR,\quad f\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2},\lambda\right)=\lambda^{2}(\lambda-q)^{2}-(M-pq)^{2},$
it follows that $\Delta(s,t;\lambda)=-\lambda^{2}t^{2}-LR+(M-pq)^{2}$. Now
$\Delta(s,t;\lambda)$ $=-\lambda^{2}t^{2}+2(M-pq)st+\left[(\lambda-q)^{2}s^{2}+(\lambda-p)^{2}t^{2}+2(\lambda-q)(\lambda-p)st\right]-s^{2}t^{2}$
$=\left((\lambda-p)^{2}-\lambda^{2}\right)t^{2}+(\lambda-q)^{2}s^{2}+2st\left[(M-pq)+(\lambda-p)(\lambda-q)\right]-s^{2}t^{2}.$
Recall that $(\lambda-p)^{2}-\lambda^{2}=p(p-2\lambda)$ and
$(M-pq)+(\lambda-p)(\lambda-q)=\lambda^{2}-(p+q-\tfrac{n}{2})\lambda=\lambda(\lambda+1),$
we obtain
$\Delta(s,t;\lambda)=p(p-2\lambda)t^{2}+(\lambda-q)^{2}s^{2}-s^{2}t^{2}+2\lambda(\lambda+1)st,$
as desired. ∎
Recall that $\xi$ is the smallest real root of the polynomial
$f\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2},\lambda\right)=\lambda^{2}(\lambda-q)^{2}-(M-pq)^{2}.$
By Lemma 2.2, for each feasible pair $(a,a_{0})$, $f(a,a_{0},\lambda)$ has exactly one real root on $(-\infty,0)$, and therefore $\xi<0$ is the unique negative root of $f\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2},\lambda\right)$ characterized by
$\lambda(\lambda-q)=-(M-pq)\quad\Longleftrightarrow\quad\lambda^{2}+(1+p)\lambda-pq=0.$ (3.5)
From (3.5) we will use the convenient identity
$\xi(\xi+1)=p(q-\xi).$ (3.6)
To prove Eq. (3.2), it suffices to prove
$\Delta(s,t;\xi)\geq 0\qquad\text{for every feasible pair }(a,a_{0}).$ (3.7)
Define
$\mu:=-\xi>0,\qquad\alpha:=p+2\mu=p-2\xi>0,\qquad\beta:=q+\mu=q-\xi>0.$
By Claim 1 and (3.6), we can rewrite
$\Delta(s,t;\xi)=(\beta^{2}-t^{2})s^{2}+2p\beta ts+p\alpha t^{2}.$
For each fixed $t$, this is a convex quadratic polynomial in $s$ with leading coefficient $\beta^{2}-t^{2}>0$ (since $\beta>q\geq|t|$). Completing the square gives
$\Delta(s,t;\xi)=(\beta^{2}-t^{2})\left(s+\frac{p\beta t}{\beta^{2}-t^{2}}\right)^{2}+\frac{pt^{2}}{\beta^{2}-t^{2}}\left[\alpha(\beta^{2}-t^{2})-p\beta^{2}\right].$ (3.8)
Hence
$\min_{|s|\leq p}\Delta(s,t;\xi)\geq\min_{s\in\mathbb{R}}\Delta(s,t;\xi)=\frac{pt^{2}}{\beta^{2}-t^{2}}\left[\alpha(\beta^{2}-t^{2})-p\beta^{2}\right].$ (3.9)
Thus, to prove Eq. (3.7), it suffices to prove
$\alpha(\beta^{2}-t^{2})\geq p\beta^{2}\qquad\text{for any }|t|\leq q.$ (3.10)
In particular, it is enough to verify that $\alpha(\beta^{2}-q^{2})-p\beta^{2}\geq 0$. Using $\mu^{2}=(1+p)\mu+pq$ from (3.5), we obtain
$\alpha(\beta^{2}-q^{2})-p\beta^{2}$ $=(p+2\mu)\mu(2q+\mu)-p(q+\mu)^{2}=2\mu^{3}+4\mu^{2}q-pq^{2}$
$=2\mu^{3}+4q((1+p)\mu+pq)-pq^{2}$
$=2\mu^{3}+4(1+p)\mu q+3pq^{2}\ >0.$
Therefore $\min_{|s|\leq p}\Delta(s,t;\xi)\geq 0$ for all $|t|\leq q$, this completes the proof of Eq. (3.1). The least root of the quadratic equation (3.6) is
$\xi=-\frac{n-\chi+2+\sqrt{(n-\chi-2)^{2}+4\chi(n-\chi)}}{4}.$
By Eq. (3.1), it follows that for every feasible pair $(a,a_{0})$, the smallest real root of $f(a,a_{0},\lambda)$ is greater than or equal to $\xi$. Hence, by Lemma 2.1, for all graphs of order $n$ with chromatic number $\chi$ satisfying $3\leq\chi\leq n-1$, the least adjacency eigenvalue $\lambda_{n}$ satisfies
$\lambda_{n}\geq\xi=-\frac{n-\chi+2+\sqrt{(n-\chi-2)^{2}+4\chi(n-\chi)}}{4}.$ (3.11)
Now we consider the equality condition. If $\lambda(G)=\xi$, then $G$ must be of the form $G(a,a_{0})$ for some feasible integers $a$ and $a_{0}$ by Lemma 2.1. Hence the smallest real root of $f\!\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2},\lambda\right)$ coincides with that of $f(a,a_{0},\lambda)$, which forces $\Delta(s,t;\xi)=0$. Since $\alpha(\beta^{2}-q^{2})-p\beta^{2}>0$, by (3.9) the only way for $\Delta(s,t;\xi)=0$ is to have $t=0$. Substituting $t=0$ into $\Delta(s,t;\xi)$ gives
$\Delta(s,0;\xi)=(\beta^{2}-0)s^{2}=\beta^{2}s^{2}.$
Thus $\Delta(s,t;\xi)=0$ implies $s=0$ as well, i.e. $(s,t)=(0,0)$. Equivalently, $a=\tfrac{\chi}{2}$ and $a_{0}=\tfrac{n-\chi}{2}$. This shows that the equality graph in (3.11) is precisely
$G\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2}\right)=\left(K_{\frac{\chi}{2}}\cup\tfrac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right),$
where both $n$ and $\chi$ are even. ∎
By the same argument as in the proof of [3, Theorem 3.2], we derive the following upper bound on the chromatic number in terms of its order and least adjacency eigenvalue.
Theorem 3.2.
Let $G$ be a graph of order $n$ with chromatic number $3\leq\chi\leq n-1$ and least adjacency eigenvalue $\lambda_{n}$. Then
$\chi\leq\left(\frac{n}{2}+1+\lambda_{n}\right)+\sqrt{\left(\frac{n}{2}+1+\lambda_{n}\right)^{2}-4(\lambda_{n}+1)\left(\lambda_{n}+\frac{n}{2}\right)},$ (3.12)
with equality if and only if $G\cong\left(K_{\frac{\chi}{2}}\cup\frac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\frac{n-\chi}{2}K_{1}\right)$, where both $n$ and $\chi$ are even.
4. A comparison with Wilf’s bound
As noted above, Wilf proved that $\chi(G)\leq 1+\lambda_{1}(G)$. In fact, as proved by Wu and Elphick [7, Lemma 3.2],
$\chi(G)\leq\chi_{\ell}(G)\leq\operatorname{col}(G)\leq 1+\lambda_{1}(G),$
where $\chi_{\ell}(G)$ denotes the list chromatic number and $\operatorname{col}(G)$ denotes the coloring number. However, in Theorem 3.2 one cannot replace $\chi(G)$ by $\chi_{\ell}(G)$ or by $\operatorname{col}(G)$.
Indeed, take $G=K_{3,3}$. Here $n=6$ and the least adjacency eigenvalue is $\lambda_{n}(G)=-3$. Plugging $\lambda=-3$ and $n=6$ into the right-hand side of (3.12) gives
$\left(\frac{n}{2}+1+\lambda\right)+\sqrt{\left(\frac{n}{2}+1+\lambda_{n}\right)^{2}-4(\lambda_{n}+1)\left(\lambda_{n}+\frac{n}{2}\right)}=2.$
But since $\operatorname{col}(K_{3,3})\geq\chi_{\ell}(K_{3,3})=3$, if $\chi$ in Theorem 3.2 were replaced by $\chi_{\ell}$ or $\operatorname{col}$, it would falsely assert that $\chi_{\ell}(K_{3,3})\leq\operatorname{col}(K_{3,3})\leq 2$. Therefore the chromatic number in Theorem 3.2 cannot be replaced by the list chromatic number or the coloring number.
Fan, Yu and Wang [3] demonstrated that their bound and Wilf’s bound are incomparable. To exemplify this incomparability, for the connected graphs in the Wolfram Mathematica database of graphs with $n=16$ and $\chi\geq 3$:
  • the Wilf bound outperforms the Fan bound for 333 graphs;
  • the Fan bound outperforms the Wilf bound for 23 graphs; and
  • the bounds are equal for 3 graphs.
The Fan–Yu–Wang bound performs comparatively well for example for some Circulant, complete multi-partite and Cone graphs.
5. Concluding Remarks
In this paper, we resolved a conjecture proposed by Fan, Yu and Wang, proving that their upper bound on the chromatic number also holds for $\frac{n}{2}<\chi\leq n-1$. It is worth noting that they also proved the following result [3, Theorem 3.3]:
Theorem 5.1 ([3]).
Among all graphs of order $n$ and with chromatic number $\chi$, where $3\leq\chi\leq n/2$, the graph $G(\lceil\chi/2\rceil,\lfloor(n-\chi)/2\rfloor)$ is the unique one whose least eigenvalue attains the minimum.
Fan, Yu and Wang [3] also conjectured that this theorem remains true for the full range $3\leq\chi\leq n-1$. This conjecture follows directly from Theorem 3.1 when both $n$ and $\chi$ are even. For the remaining parity cases, our fully algebraic approach appears to be less straightforward than the calculus-based approach used in [3]. At the symmetric point $(a,a_{0})=(\tfrac{\chi}{2},\tfrac{n-\chi}{2})$, the characteristic polynomial can be factorized as a difference of squares, and thus the least root $\xi$ satisfies a quadratic equation. Substituting this relation into $\Delta$ is straightforward, and the nonnegativity of $\Delta$ can then be easily verified. However, for instance, when both $\chi$ and $n$ are odd, we have $(a,a_{0})=(\tfrac{\chi+1}{2},\tfrac{n-\chi}{2})$. In this case, the symmetry is broken, and the least root $\zeta$ instead satisfies a quartic equation:
$\lambda^{2}(\lambda-q)^{2}-\left(\frac{n}{2}\lambda-pq\right)^{2}+\tfrac{1}{4}p(p-2\lambda)=0.$
Consequently, proving $\tilde{\Delta}(s,t;\zeta):=f(a,a_{0},\zeta)-f\left(\tfrac{\chi+1}{2},\tfrac{n-\chi}{2},\zeta\right)\geq 0$ must be carried out over the half-integer lattice $t\in\tfrac{1}{2}+\mathbb{Z}$, with separate treatment required for the endpoint cases $t=\tfrac{1}{2},\tfrac{3}{2},$ and $t=q$ (since we may assume without loss of generality that $a\geq b$, i.e., $t\geq 0$). Each of these endpoint cases involves repeated eliminations using the above quartic identity, which is substantially more complicated than the algebraic substitution based on a quadratic equation. We believe the conclusion still holds, although a concise and unified algebraic proof covering all parity cases remains elusive.
Acknowledgements
The authors would like to thank Prof. Jie Ma for helpful comments and suggestions.